The temperature of a metal having 7 g mass was raised by 117⁰C after 237.5 J of energy was supplied. Determine what is the metal's specific heat.


1
Expert's answer
2022-01-20T10:08:39-0500

Solution

Energy is given

"\Delta Q=237.5J"

Mass of metal M=0.007kg

Increasement in temperature

T=117°C=380K

So specific heat is defined as below formula

"S=\frac{\Delta Q}{mT}\\=\frac{237.8J}{0.007kg\times380K}\\=89.40\frac{J}{kg-K}"